Tech Staff Meeting Notes March 6, 2017

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

Fish Fry will be Friday May 5, Cinco de Mayo. Cost remains $6.

Arch asked about updating the signpost outside the middle house.


Hardware for ET network is entered. John and James will work together to load dataless.

Dave and James have started loading hardware for NM and AG.


GX620 at mttnworm needs to be visually checked. All others with "puffy caps" have been repaired. We have five spare motherboards.


Data taken from: /gaia/smeagol/dat2/Working/NM/201703/20170303_202533_MAN

Misc Problems
BACM- spikes
BRNM - "
BOAR - "
CPAR - "
LVAR - "
TMAR - "
TYAR - "
LPAR - dropping out
HBAR - "

Low solar dropouts have several different sources though common problem is an extraordinarily overcast January. Other factors include mis-wiring, ISIS3 xmiters draw more power (so need more solar), AGM's have problems with low power charge, etc. Jim's spreadsheet shows 9 stations that need an upgrade, 10 on the edge, and 13 that need to be checked.

Need a site for MORT2.

There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.

I40 Bridge comms.
We're testing a pair of Afar radios from Jim Smith. They are promising. We'd have two access points at the Law School connected to a pair on the bridge. The DAS's on the bridge would communicate with each other under the bridge. The access points automatically balance the load between the two.

The spectre routers have a relatively short character limit for url's. But when you change a config using the web interface with the obsidians, they try to resend the entire configuration using a url and its gets truncated when coupled with a spectre router. So make sure the configs are correct because it requires a site visit to change them.

NODE Tasks

All nodes are due a firmware upgrade that will sufficiently change things such that the the initial upgrade and configure should be done at CERI, then swap hardware at the node. This will be done as machines are replaced.

DBB: Rebuild SFTN-MKTA portion of DBB. Done. Upgrade complete to Luxora. Hardware is on hand to upgrade the rest of the DBB. The links from Luxora to Edith and Edith to Lennox are next up and will be done at simultaneously. Be sure to schedule with James so he can temporarily reroute data traffic to the dsl.


Property at MKTA node will possibly be sold. Chris is looking into to it to make sure we can retain our presence.

Getting interference at the node causing spikes in CPAR, LVAR, LRAR, and TYAR.



AZTN/HDBR: Need to pull the 200' sensor at the weigh station. Probable water damage.


We have no dbb firewall spare. Steve will get a pair of replacements.

Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave. There are several open work orders. Jim will make a prioritized list and speak with zone supervisor.


Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob (MaggieValley), a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Compressor noise suppression is scheduled for the end of the month afterwhich a new noise study can be done.


Data taken from: /gaia/smeagol/dat2/Working/ET/201702/20170226_022300_MAN
It's been very quiet so I'm not sure if this is the entire list of problems.


Misc Problems

GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.

John's updated todo list .

Spikes at CVVA most likely originating at flat gap.


Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.


HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Competion of fiber and supporting hardware should be done, we're told, by late April. Need to add the node end of the radio link for GFM2.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM2 is good to go. It will be a reftek with S-13's. Do we need to take student volunteers to help with the GFM decommission?


WMTN pit needs to be moved due to erosion.

CPRT comm reconfig. John says, "CPRT is going through the Senior Center at Corryton (cort). We removed everything from svfd an everything but what is on the tower at gttn. cort is indeed using dynamic addressing but data is flowing. Greg Hill the director at the senior center is supposed to be contacting his IT person to see what can be done. I left a message for Greg last week but still haven't heard back. I want to get Greg's IT guy in direct contact with either Steve or James so they can figure it all out."


Site for newstation is permitted at a school in Armorel, AR just east of Blytheville. The monument is installed and needs to be welded. The host wants a brown privacy fence ASAP. The PI is making things difficult by cutting corners and trying to do everything on the cheap.

CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.

AG Network

X40A upgrade to AG standard is complete. Still have U40A and Z41A to do.

Leak at FCAR is repaired.

When the water leak at LCAR was repaired, was an accelerometer reinstalled? Yes.


Needs some clean up.

The FDL wiki is here:


The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.